Grade 11Physical Chemistry

For the decomposition reaction of CaCO3,the Kp=1.16 atm at 8000c
If 20g of CaCO3 were kept in a 10 lit container and heated upto 8000c,what percentage of CaCO3 would remain unreacted at equilibrium

vinodh
11 Years agoGrade 11

2 Answers

Sunil Kumar FP
11 Years ago
molecular mass of CaCO3=100
mole of CaCO3=20/100 =.2 mole
molarit of CaCO3=.2/10 =.02Lt
CaCO3 ------------->CaO +CO2
.02 0 0
.02-x x x
Kp =pCO2
1.16=x*.082*1073
x=.013
unreacted=.007
percentage=35%
Tahir hussain
7 Years ago
The value for caco3--´`gives cao+ co2 at 727ہCis 8.21atm.If 100g  of caco3 was heated in a 3L closed vessel at 273ہC.What percent of sample with remain unreacted at equilibrum.